OpenStudy (anonymous):

Tom has 80 nickels and dimes $5.15. Write and solve an equation to determine how many of each coin he has. Explain

OpenStudy (anonymous):

\[10x+5(80-x)=515\]

OpenStudy (anonymous):

oh explanation x is number of dimes so 80-x is number of nickles. dimes are worth ten cents and nickles 5 and the total is 515 (cents) so that is the equation

OpenStudy (anonymous):

so how many of each coin do you have then... can you show me the steps on how to solve it

OpenStudy (anonymous):

STEPS

OpenStudy (anonymous):

damn who ever gave you this problem slap them

OpenStudy (anonymous):

HAHA can u help me

OpenStudy (anonymous):

i am sry buh it just gt me lost by lookin at it

OpenStudy (radar):

Continue with what satellite73 has provided: x= the number of dimes 80-x = the number of nickels 10x +5(80-x)=515 Now work that out, I will wait to see how you do.

OpenStudy (anonymous):

radar looks smart can u solve this out too :What does the consumer price index measure?

OpenStudy (anonymous):

not to be rude between you two buh i needed help also man

OpenStudy (anonymous):

radar help me solve the prob

OpenStudy (radar):

The Consumer Price Index (CPI) is a measure of the average change over time in the prices paid by urban consumers for a market basket of consumer goods and services. I do not know what comprises this market basket and services. Google it and you will find it.

OpenStudy (anonymous):

radr can you help me next if yu can buh help bailey first ....thks

OpenStudy (radar):

Can you work equations? Do you know how to work with parenthesis?

OpenStudy (anonymous):

thanks i got my answer

OpenStudy (anonymous):

just show me the steps

OpenStudy (radar):

I would rather you show me the steps you would take to solve this equation: 10x +5(80-x)=515 do the multiplication first, do the 5(80-x)

OpenStudy (radar):

If you can not do the simple things, then you need to review. For this case BaileyC review the distributive properties.

OpenStudy (anonymous):

i did that now , can u help with the next stop

OpenStudy (radar):

What did you get for the 5(80-x) ?

OpenStudy (anonymous):

400-5x

OpenStudy (radar):

Excellent, now combine that with the 10x 10x+400-5x

OpenStudy (anonymous):

just giver her the damn answer man make life easy

OpenStudy (anonymous):

i solved it bye

OpenStudy (radar):

easy for who mishi? Do you think going through life needing help is easy?

OpenStudy (anonymous):

easy for that girl who was askin you nad yea in life needin help is easy if you ask the right way with pateience

OpenStudy (radar):

Haven't you heard, the old saying, Give a fish to a hungry man, and the next day he will be hungry again, teach a man how to fish, he won't be hungry again.

OpenStudy (anonymous):

well your right and no it was my first time hearing this saying buh honestly it hitted me on my head ...i am sorry

OpenStudy (anonymous):

ok is the answer 82 dimes and 72 nickles

OpenStudy (radar):

How much money is 82 dimes?

OpenStudy (anonymous):

?? can u just help me

OpenStudy (anonymous):

show me step by stp

OpenStudy (radar):

I was trying to but you refuse to make any effort to solve the thing. once again can you combine the like terms in this expression? 10x+400-5x

OpenStudy (anonymous):

i got that 15x+400=515

OpenStudy (radar):

10x-5x when combined is not 15x. Do you know why?

OpenStudy (anonymous):

yeah

OpenStudy (radar):

Why

OpenStudy (radar):

You didn't notice the sign was negative for the 5x. If you are going to be oblivious to the obvious then you need to regroup and focus. Bye and good luck, you are going to need it.

OpenStudy (anonymous):

yea good uck

OpenStudy (anonymous):

radar think positiv

OpenStudy (radar):

I try. .lol